# A coupled-channel perspective analysis on bottom-strange molecular pentaquarks

### v1

@@ III.1 Bottom baryon and anti-strange meson systems
Besides, we have $I(J^{P})=1/2(3/2^{-})\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ channel for $Y_{b}$ and $\bar{K}^{(*)}$ system, and the corresponding results are listed in Table 3. According to our estimations, a bound state exists below $\Lambda_{b}\bar{K}^{*}$ threshold, dominated by $\Lambda_{b}\bar{K}^{*}(^{4}D_{3/2})$ and $\Sigma_{b}\bar{K}^{*}(^{4}S_{3/2})$ channel, and is sensitive to the cutoff value. Since the cutoff value consists with the empirical value for deuteron, the $I(J^{P})=1/2(3/2^{-})\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ can be regarded as a possible molecular state candidate. In the $I(J^{P})=3/2(3/2^{-})$ $\Sigma_{b}\bar{K}^{*}$ system, a weakly bound state with energy of $6703$ $\rm{MeV}$ appears at cutoff $1300$ $\rm{MeV}$ and is dominated by $S-$wave channel. If only the one pion exchange potential is considered, a bound state is obtain with cutoff $2400$ $\rm{MeV}$, which means that the potentials of $\rho$ and $\omega$ exchanges are helpful to form the bound state.
@@ III.2 Bottom baryon and strange meson systems
As for the $\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system with $I(J^{P})=1/2(3/2^{-})$, at the cutoff is 1260 $\rm MeV$, a bound state below the $\Lambda_{b}K^{*}$ threshold emerges. One can also find that when only $\Sigma_{b}K^{*}$ channel is considered, a loosely bound state appears, which is listed in Table 3. Finally, for the $I(J^{P})=3/2(3/2^{-})$ $\Sigma_{b}K^{*}$ system, we can not obtain any bound state solution with $\Lambda=800\sim 5000$ MeV.
Table 4: The summary of our predictions for bottom strange pentaquark molecular state systems with cutoff $\Lambda$ in a range of $800\sim 1100$ $\rm MeV$. Here, the , ""("×")""""$\checkmark$"($\times$")" ✓ " ( " × " ) represents that the corresponding state may (may not) form a molecular state.
$I(J^{PC})$ Mass$(\rm MeV)$ Width$(\rm MeV)$ $r_{RMS}(\rm fm)$ Status Selected decay mode
12($\frac{1}{2}(\frac{1}{2}^{-})$$\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ $6301\sim 6222$- $2.49\sim 0.73$checkmark $\Lambda_{b}\bar{K}/\Xi_{b}^{(\prime)}\pi$
$6693\sim 6516$ $34.58\sim 31.00$ $1.70+1.43i\sim 0.61+0.51i$checkmark $\Lambda_{b}\bar{K}^{(*)}/\Sigma_{b}\bar{K}/\Lambda\bar{B}^{(*)}/\Sigma\bar{B}/\Xi_{b}^{(\prime)}\pi/\Xi_{b}^{(\prime)}\eta/\Xi_{b}\rho/\Xi_{b}\omega$
$\frac{3}{2}(\frac{1}{2}^{-})$$\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$- -- -- $\times$ --
$\frac{1}{2}(\frac{3}{2}^{-})$$\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$- -- -- $\times$ --
$\frac{3}{2}(\frac{3}{2}^{-})\Sigma_{b}\bar{K}^{*}$ $6703\sim 6702$- $3.37\sim 2.62$checkmark $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}^{*}\bar{K}/\Lambda\bar{B}^{*}/\Sigma\bar{B}^{*}/\Xi_{b}^{*}\pi/\Xi_{b}^{*}\eta/\Xi_{b}\rho/\Xi_{b}\omega$
12($\frac{1}{2}(\frac{1}{2}^{-})$$\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ $6303\sim 6269$- $3.13\sim 1.08$checkmark $N\bar{B}_{s}/\Lambda_{b}K$
$\frac{3}{2}(\frac{1}{2}^{-})$$\Sigma_{b}K/\Sigma_{b}K^{*}$- -- -- $\times$ --
$\frac{1}{2}(\frac{3}{2}^{-})$$\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$- -- -- $\times$ --
32($\frac{3}{2}(\frac{3}{2}^{-})$$\Sigma_{b}K^{*}$ $6704\sim 6692$- $4.13\sim 1.34$checkmark $N\bar{B}_{s}^{*}/\Lambda_{b}^{*}K/\Sigma_{b}^{*}K$
@@ III.3 Further discussions
For $Y_{b}\bar{K}^{(*)}$ systems, we can obtain bound states and resonances, but only bound states are revealed for $Y_{b}K^{(*)}$ systems. The reason is that the flavor factors in the potentials for these systems are quite different, which determine the relative sign and strength and are crucial for the formation of molecular states. It is also interesting to note that the root mean square (RMS) radius $r_{RMS}$ for the $I(J^{PC})=1/2(1/2^{-})\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ resonances can be a complex number. In such cases, one can use the interpretation scheme proposed by T. Berggren, which generalizes the concept of expectation values from bound states to resonances Berggren:1970wto . According to this scheme, the real part of the complex $r_{RMS}$ represents the usual physical expectation value, while the imaginary part indicates a measure of uncertainty in the observation. Numerical calculations of $r^{2}$ have supported this generalized interpretation  Gyarmati:1972yac ; 1997matrix . We illustrate the The complex energy eigenvalues of $I(J^{P})=1/2(1/2^{-})$ system with varying the angle $\theta$ from $30^{\circ}\sim 40^{\circ}$ in Figure 3.
According to the masses and quantum numbers, we present some possible decay channels for these predicted states in Table 4. For instance, the $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ bound state can be found in $\Lambda_{b}\bar{K}$ and $\Xi_{b}^{(\prime)}\pi$ channels. In the literature Chen:2018orb ; He:2021xrh ; Wang:2018fjm ; Cui:2019dzj , both $J^{P}=\frac{1}{2}^{-}$ molecular and $J^{P}=\frac{3}{2}^{-}/\frac{5}{2}^{-}$ conventional interpretations exist for the particle $\Xi_{b}(6227)$, and the spin is certainly crucial for distinguishing these two explanations. Another way to solve this puzzle is to hunt for the flavor exotic state with $6303\sim 6269~{}\rm{MeV}$ in the $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system, which is the mirror state of $\Xi_{b}(6227)$ in the molecular picture but does not appear in the three-quark picture. We highly hope that the future experiments can verify our proposals.
@@ IV Summary
In this work, we systematically investigate the coupled $Y_{b}\bar{K}^{(*)}(Y_{b}K^{(*)})$ system to search for possible bound states and resonances by adopting one-boson-exchange model within complex scaling method. For the coupled $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ systems, according to our estimations, a bound state solution is obtained, which may correspond to the observed particle $\Xi_{b}(6227)$. Meanwhile, we find a $I(J^{PC})=1/2(1/2^{-})$ resonance near the $\Sigma_{b}\bar{K}^{*}$ threshold and a bound state in the $I(J^{P})=1/2(3/2^{-})$ $\Sigma_{b}\bar{K}^{*}$ system.
Then, when we extend our study to the $Y_{b}K^{(*)}$ systems, two loosely bound states are obtained. It is worth pointing out that the predicted bound state with $6303\sim 6269~{}\rm{MeV}$ in the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system is flavor exotic and does not appear in the spectroscopy of conventional baryons, which provides a practical way to resolve the puzzle of particle $\Xi_{b}(6227)$. We hope our predictions can offer valuable information to the future experiments observations.
@@ ACKNOWLEDGMENTS
We would like to thank Rui Chen, Xian-Hui Zhong, Li-Cheng Sheng, and Jin-Yu Huo for useful discussions. The work of X.-N. X. and Q. F. Song is supported by the National Natural Science Foundation of China under Grants No. 12275364. Q.-F. Lü is supported by the Natural Science Foundation of Hunan Province under Grant No. 2023JJ40421, the Scientific Research Projects of Hunan Provincial Education Department under Grant No. 24B0063, and the Youth Talent Support Program of Hunan Normal University under Grant No. 2024QNTJ14.

### v2

@@ III.1.4 $I(J^{P})=3/2(3/2^{-})$ with $\Sigma_{b}\bar{K}^{*}$
In the $I(J^{P})=3/2(3/2^{-})$ $\Sigma_{b}\bar{K}^{*}$ system, a weakly bound state with an energy of $6703$ $\rm{MeV}$ appears at a cutoff of $1300$ $\rm{MeV}$ and is dominated by the $S$-wave channel about 99%. If only the one-pion-exchange potential is considered, a bound state is obtained with a cutoff of $2400$ $\rm{MeV}$, which means that the potentials of $\rho$ and $\omega$ exchanges are helpful to form the bound state.
@@ III.2 Bottom baryon and strange meson systems
For $Y_{b}K^{(*)}$ systems, the effective potentials from the $\omega$ and $\pi$ exchanges are in complete contrast with those of the $Y_{b}\bar{K}^{(*)}$ systems. Unlike the bottom baryon and anti-strange meson systems, for the bottom baryon and strange meson systems, one can only obtain bound state solutions. The corresponding numerical results are collected in Table 3 and Figure 3(b).
III.2.1 $I(J^{P})=1/2(1/2^{-})$ with $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$
When the cutoff parameter varies within $920\sim 1100~\rm{MeV}$, the mass decreases from 6304 to 6269 $\rm{MeV}$, while the root-mean-square radius $r_{RMS}$ shrinks from 3 to 1 $\rm{fm}$, consistent with expectations for a molecular candidate. Considering only the single $\Sigma_{b}K$ channel, a bound state solution also emerges at a larger cutoff of $\Lambda=1900~\rm{MeV}$, further indicating that coupled-channel effects play a significant role in forming the molecule.
III.2.2 $I(J^{P})=3/2(1/2^{-})$ with $\Sigma_{b}K/\Sigma_{b}K^{*}$
For the $I(J^{P})=3/2(1/2^{-})$ $\Sigma_{b}K/\Sigma_{b}K^{*}$ system, a bound state solution is found when the cutoff lies in the $1260~\rm MeV$, which is mainly composed of $S$- wave $\Sigma_{b}K$ and $\Sigma_{b}K^{*}$ about $76\%$ and 24%, respectively. When we vary the cutoff to $1270~\rm MeV$, the predicted mass varies from 6298 to 6284 $\rm{MeV}$, with the corresponding $r_{RMS}$ decreasing from $1.5$ to $0.9~\rm{fm}$. This sensitivity to the cutoff suggests the possibility of a molecular state.
III.2.3 $I(J^{P})=1/2(3/2^{-})$ with $\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$
As listed in Table 3, for the $\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system with $I(J^{P})=1/2(3/2^{-})$, a bound state below the $\Lambda_{b}K^{*}$ threshold appears at a cutoff of 1260 $\rm{MeV}$. However, if only the $\Lambda_{b}K^{*}$ channel is considered, no bound state pole is found for cutoff values ranging from 800 to 5000 $\rm MeV$. In contrast, a loosely bound state dominated by the $\Sigma_{b}K^{*}(^{4}S_{3/2})$ component (about 98%) appears when considering only the $\Sigma_{b}K^{*}$ channel, with a cutoff of 960 $\rm MeV$. This state has a mass of 6704 $\rm MeV$ and a root-mean-square radius $r_{\rm RMS}=4~\rm fm$.
III.2.4 $I(J^{P})=3/2(3/2^{-})$ with $\Sigma_{b}K^{*}$
Finally, for the $I(J^{P})=3/2(3/2^{-})$ $\Sigma_{b}K^{*}$ system, no bound state solution is found within the cutoff range $\Lambda=800\sim 5000~\rm{MeV}$.
@@ III.2.4 $I(J^{P})=3/2(3/2^{-})$ with $\Sigma_{b}K^{*}$
[…]
$\frac{1}{2}(\frac{1}{2}^{-})$$\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ $6301\sim 6222$ - $2.49\sim 0.73$ $\checkmark$ $\Lambda_{b}\bar{K}/\Xi_{b}^{(\prime)}\pi$
$6693\sim 6516$ $34.58\sim 31.00$ $1.22-0.55i\sim 0.84-0.10i$ $\checkmark$ $\Lambda_{b}\bar{K}^{(*)}/\Sigma_{b}\bar{K}/\Lambda\bar{B}^{(*)}/\Sigma\bar{B}/\Xi_{b}^{(\prime)}\pi/\Xi_{b}^{(\prime)}\eta/\Xi_{b}\rho/\Xi_{b}\omega$
$\frac{3}{2}(\frac{1}{2}^{-})$$\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$ - - - $\times$ -
$\frac{1}{2}(\frac{3}{2}^{-})$$\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ - - - $\times$ -
$\frac{3}{2}(\frac{3}{2}^{-})\Sigma_{b}\bar{K}^{*}$ $6703\sim 6702$ - $3.37\sim 2.62$ $\checkmark$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}^{*}\bar{K}/\Lambda\bar{B}^{*}/\Sigma\bar{B}^{*}/\Xi_{b}^{*}\pi/\Xi_{b}^{*}\eta/\Xi_{b}\rho/\Xi_{b}\omega$
$\frac{1}{2}(\frac{1}{2}^{-})$$\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ $6303\sim 6269$ - $3.13\sim 1.08$ $\checkmark$ $N\bar{B}_{s}/\Lambda_{b}K$
$\frac{3}{2}(\frac{1}{2}^{-})$$\Sigma_{b}K/\Sigma_{b}K^{*}$ - - - $\times$ -
$\frac{1}{2}(\frac{3}{2}^{-})$$\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ - - - $\times$ -
$\frac{3}{2}(\frac{3}{2}^{-})$$\Sigma_{b}K^{*}$ $6704\sim 6692$ - $4.13\sim 1.34$ $\checkmark$ $N\bar{B}_{s}^{*}/\Lambda_{b}^{*}K/\Sigma_{b}^{*}K$
@@ III.3 Further discussions
For $Y_{b}\bar{K}^{(*)}$ systems, we can obtain bound states and resonances, but only bound states are revealed for $Y_{b}K^{(*)}$ systems. The reason is that the flavor factors in the potentials for these systems differ significantly, which determine the relative sign and strength and are crucial for the formation of molecular states.
In our study, we aim to explore the crucial role of coupled-channel effects in the production of resonant states. As shown in Table 4, when considering only the single-channel $\Sigma_{b}\bar{K}^{*}$ $S$-$D$ wave mixing, we obtain a bound state at a cutoff of 840 $\rm{MeV}$. However, when including the coupled-channel effect for the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$ system, we find a relatively narrow resonant state at a cutoff of 900 $\rm{MeV}$, which is dominated by the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}^{*}(^{2}S_{1/2})$ component. Additionally, for the coupled-channel system $I(J^{P})=1/2(1/2^{-})$ $\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$, we obtain a relatively broad resonance at a cutoff of 920 $\rm{MeV}$. Furthermore, when considering the fully coupled $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ system, a resonance emerges near the $\Sigma_{b}\bar{K}^{*}$ threshold, with a complex energy $E{-}i\Gamma/2=6692.84{-}17.29i$ at a cutoff of 960 $\rm{MeV}$. On the other hand, when we discard the $D$-wave contributions, we still obtain a resonant state at a cutoff of approximately 1000 $\rm{MeV}$ for the $S$-wave coupled systems $\Sigma_{b}\bar{K}/\Sigma_{b}\bar{K}^{*}$, $\Lambda_{b}\bar{K}{/}\Sigma_{b}\bar{K}^{*}$, and $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ with $I(J^{P})=1/2(1/2^{-})$, respectively. In conclusion, our calculations demonstrate that the coupled-channel effect plays a significant role in the formation of resonant states, with the main contributions coming from the $S$-wave.
According to the masses and quantum numbers, we present some possible decay channels for these predicted states in Table 5. For instance, the $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ bound state can be found in the $\Lambda_{b}\bar{K}$ and $\Xi_{b}^{(\prime)}\pi$ channels. In the literature Chen:2018orb ; He:2021xrh ; Wang:2018fjm ; Cui:2019dzj , both $J^{P}=\frac{1}{2}^{-}$ molecular and $J^{P}=\frac{3}{2}^{-}/\frac{5}{2}^{-}$ conventional interpretations exist for the particle $\Xi_{b}(6227)$, and the spin is certainly crucial for distinguishing these two explanations. Another way to solve this puzzle is to hunt for the flavor exotic state with mass $6303\sim 6269~\rm{MeV}$ in the $I(J^{PC})=1/2(1/2^{-})$ $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system, which is the mirror state of $\Xi_{b}(6227)$ in the molecular picture but does not appear in the three-quark picture. We strongly hope that future experiments can verify our proposals.
@@ IV Summary
In this work, we systematically investigate the coupled $Y_{b}\bar{K}^{(*)}(Y_{b}K^{(*)})$ system to search for possible bound states and resonances by adopting the one-boson-exchange model within the complex scaling method. For the coupled $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}\bar{K}/\Lambda_{b}\bar{K}^{*}/\Sigma_{b}\bar{K}^{*}$ systems, according to our estimations, a bound state solution is obtained, which may correspond to the observed particle $\Xi_{b}(6227)$. Meanwhile, we find a $I(J^{PC})=1/2(1/2^{-})$ resonance near the $\Sigma_{b}\bar{K}^{*}$ threshold and a bound state in the $I(J^{P})=1/2(3/2^{-})$ $\Sigma_{b}\bar{K}^{*}$ system.
Then, when we extend our study to the $Y_{b}K^{(*)}$ systems, two loosely bound states are obtained. It is worth pointing out that the predicted bound state with mass $6303\sim 6269~\rm{MeV}$ in the $I(J^{P})=1/2(1/2^{-})$ $\Sigma_{b}K/\Lambda_{b}K^{*}/\Sigma_{b}K^{*}$ system is flavor exotic and does not appear in the spectroscopy of conventional baryons, which provides a practical way to resolve the puzzle of the particle $\Xi_{b}(6227)$. We hope our predictions can offer valuable information to future experimental observations.
@@ ACKNOWLEDGMENTS
We would like to thank Rui Chen, Xian-Hui Zhong, Li-Cheng Sheng, and Jin-Yu Huo for useful discussions. The work of X.-N. X. and Q. F. Song is supported by the National Natural Science Foundation of China under Grants No. 12275364. Q.-F. Lü is supported by the Natural Science Foundation of Hunan Province under Grant No. 2023JJ40421, the Scientific Research Foundation of Hunan Provincial Education Department under Grant No. 24B0063, and the Youth Talent Support Program of Hunan Normal University under Grant No. 2024QNTJ14.